OpenStudy (anonymous):

Simplify each number using i. I've had this explained so many times and I just don't get it..

OpenStudy (anonymous):

sqrt -7 sqrt -15 sqrt -32 sqrt -72

jimthompson5910 (jim_thompson5910):

It turns out that \[\Large i = \sqrt{-1}\] so what we do is factor each radicand (the stuff under the root) and break up the root like so \[\Large \sqrt{-7} = \sqrt{-1*7}\] \[\Large \sqrt{-7} = \sqrt{-1}*\sqrt{7}\] \[\Large \sqrt{-7} = i\sqrt{7}\]

jimthompson5910 (jim_thompson5910):

The second one is done the exact same way \[\Large \sqrt{-15} = \sqrt{-1*15}\] \[\Large \sqrt{-15} = \sqrt{-1}*\sqrt{15}\] \[\Large \sqrt{-15} = i\sqrt{15}\]

jimthompson5910 (jim_thompson5910):

The third one is a bit different. You need to factor out the -1 like before, but you also have to factor out 16 since this is the largest perfect square factor. \[\Large \sqrt{-32} = \sqrt{-1*16*2}\] \[\Large \sqrt{-32} = \sqrt{-1}*\sqrt{16}*\sqrt{2}\] \[\Large \sqrt{-32} = i*4*\sqrt{2}\] \[\Large \sqrt{-32} = 4i\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

I'll let you do the last one. Tell me what you get.

OpenStudy (anonymous):

thanks i'll do it right now

OpenStudy (anonymous):

6i2−−√

OpenStudy (anonymous):

6isqrt2

jimthompson5910 (jim_thompson5910):

Very good, \[\Large \sqrt{-72} = 6i\sqrt{2}\] is correct

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

you're welcome